OpenStudy (anonymous):

What rule represents the function. (0,0)(1,0.5)(2,2)(3,4.5)

OpenStudy (anonymous):

Please help

OpenStudy (aum):

x y factor (that is, x times ? = y) 0 0 0 1 0.5 0.5 2 2 1.0 3 4.5 1.5 ..... x y y = x * x / 2 = x^2/2 y = x^2 / 2 is the rule.

OpenStudy (anonymous):

Oh my gosh thank you thank you thank you

OpenStudy (aum):

You are welcome.